A university wants to examine the effect of financial aid on graduation rate. Among the students, 90% of them graduate with a bachelor degree. For those who graduate, 75% of them receive financial aid. For those who do not graduate, 50% of them receive financial aid. For those who receive financial aid, what is the probability that they graduate?
Answer #1

P(receive financial aid) = 0.75 * 0.9 + 0.5 * (1 - 0.9) = 0.725

P(graduate | receive financial aid) = P(receive financial aid | graduate) * P(graduate)/P(receive financial aid)

                                                      = (0.75 * 0.9)/0.725 = 0.93103